OpenStudy (anonymous):

How many atoms of nickel can be found in a 1.00 g sample? somone help pleaseeeeeeeeeeeeeeeeeeeeeeeeeeewith a cherry on top c:

OpenStudy (aaronq):

use n=m/MW ..moles equals mass divided by molecular weight n= (1g)/(58.69g/mol) = since 1 mole = 6.023x10^23 atoms (avogadros number) multiply the amount of moles you have by avogadros number and voila

OpenStudy (anonymous):

okay whatttttttttttttt? I dont understand the question at all...or your answer please help! <3

OpenStudy (aaronq):

do you know what a mole is?

OpenStudy (aaronq):

so you have 1 g of nickel right? a formula you should know is n=m/MW, which is moles equals mass divided by molecular weight (or more accurately, mass). using that formula you can find the number of moles, then to find the number of ATOMS, you multiple however many moles you have by avogadros number (6.023x^23).

OpenStudy (aaronq):

do you understand

OpenStudy (anonymous):

we just learned this in class today , the teacher is mean -.-.....I kinda understand but I'm pretty sure I'mma get this wrong anyway. Thanks though!! :)

OpenStudy (aaronq):

you go to school on sunday? lol well know what a mole is a proportion like saying a dozen. 1 dozen of atoms = 12 atoms 1 mole of atoms = 6.023x10^23 atoms use n=m/MW and it's basically plug and chug

OpenStudy (anonymous):

Its monday here in florida , lol. where are YOU from ? hmmm , I'm gonna try what you say , brb. :D

OpenStudy (aaronq):

LOL oh yeah, I'm in canada and it's thanksgiving. do it, let me know what you get to make sure its correct !

OpenStudy (anonymous):

14.56......I'm not sure If I did this right :/

OpenStudy (anonymous):

lol... Happy thanksgiving!

OpenStudy (aaronq):

thanks! haha okay so you divided (1g)/(58.69g/mol)= 0.017 moles right? now you multiply 0.017 x (6.023x^10^23) (make sure you know how to input \[6.023x10^{23}\] into your calc) you should get 1.026 x 10^22

OpenStudy (anonymous):

wth , I don't understand this at all ...this is my worst subject ... THANKS SO MUCH THOUGH! why arent you celebrating thanksgiving? lol

OpenStudy (aaronq):

lol what don't you understand? haha il try to explain i already did, had turkey n stuffing and pumpkin pie. now i'm working on a lab, due to trmw but i'm stuck :S

OpenStudy (anonymous):

stuck on what ? and I'm okay, you've helped enough! :) and yummy... I had cereal for lunch -.- lol

OpenStudy (anonymous):

what grade are you in ?

OpenStudy (aaronq):

stuck on the interpretation of nitro- nitrite- isomers on the infrared spectrum.. haha I'm in uni 2nd year.

OpenStudy (aaronq):

cereal does not cut it haha

OpenStudy (anonymous):



OpenStudy (anonymous):

lol....I'm in like 10th grade

OpenStudy (aaronq):

haha enjoy school while it's easy